# Label Smoothing for Enhanced Text Sentiment Classification

### 11 Dec 2023 v1

@@ 4 Experimental Results
We construct five LS prediction models incorporating varying degrees of LS and assess their performance compared to a baseline model on eight data sets using both BERT, TEXTCNN and RoBERTa architectures. The experiments are conducted to evaluate the effectiveness of LS methods in improving the prediction accuracy and robustness of the models.
4.1 Data Sets
@@
We utilize publicly available data sets obtained from Kaggle and Hugging Face for our study. The data set collection comprises eight distinct data sets, of which six are three-class and two are two-class. The diverse range of data sets enables us to evaluate the performance and generalizability of our models across various data types and classification scenarios.
@@
Twitter Financial News Sentiment(TFNS) was collected using the Twitter API. It holds 11,932 documents annotated with 3 labels and comprises annotated finance-related tweets.
@@
Kaggle Financial Sentiment(KFS) is a dataset of financial reviews annotated with 3 labels. These reviews cover topics such as stocks, investments, market analysis, corporate performance, technology, and real estate.
@@
Tweet Sentiment Extraction(TSE) is a three-class dataset from a Kaggle contest. It has 27481 samples of a tweet and a sentiment label.
@@
Auditor Sentiment(AS) gathers the auditor evaluations into one dataset. It contains thousands of sentences from English financial news, grouped into three categories by emotion.
@@
Financial Phrasebank(FP) consists of 4840 sentences from English language financial news categorised by sentiment. The dataset is divided into three classes by agreement rate of 5-8 annotators.
@@
Chatgpt Sentiment Analysis(CSA) contains 10,000 pieces of data. The dataset gathers tweets about Chatgpt and Reflects people’s perception of it. Emotions are divided into three categories.
@@
Rotten Tomatoes Reviews(RTR) consists of 10,662 processed sentences from Rotten Tomatoes movie reviews. It’s a two-class dataset with 5,331 positive and 5,331 negative.
@@
Sentiment140(Sent140) is a widely used sentiment analysis dataset created by researchers at Stanford University. It contains 1,600,000 tweets extracted using the twitter api. The tweets have been annotated to two classes. In this work, we only use parts of the data.
4.2 Models and Loss Functions
@@
Models: Table 1 presents the configuration of the models and loss functions used in the experiment. The LS models (LS1-LS5) employ varying levels of LS, with the smoothing intensity increasing from LS1 to LS5. The original labels are used as they are for LS1, while for LS2 to LS5, the labels are smoothed by adding small perturbations. By gradually increasing the smoothing level from LS1 to LS5, we explore the effects of different degrees of LS on the model’s accuracy and stability. All LS models utilize the KL Loss as the loss function. This loss function measures the divergence between the predicted probability distribution and the smoothed label distribution. In contrast, the baseline model uses single labels without any smoothing and employs the Cross Entropy loss function.
@@
Table 1: Model smoothing levels and loss functions. The original one-hot label is $\mathbf{y}$, and the soft label $\mathbf{y}^{LS}$
@@
Model Smoothing Level (three-class) Smoothing Level (two-class) Loss Function
LS2 $\mathbf{y}^{LS}=\mathbf{y}*0.97+0.01*\mathbf{1}$ $\mathbf{y}^{LS}=\mathbf{y}*0.98+0.01*\mathbf{1}$ KL Loss
LS3 $\mathbf{y}^{LS}=\mathbf{y}*0.925+0.025*\mathbf{1}$ $\mathbf{y}^{LS}=\mathbf{y}*0.9+0.05*\mathbf{1}$ KL Loss
LS4 $\mathbf{y}^{LS}=\mathbf{y}*0.85+0.05*\mathbf{1}$ $\mathbf{y}^{LS}=\mathbf{y}*0.8+0.1*\mathbf{1}$ KL Loss
LS5 $\mathbf{y}^{LS}=\mathbf{y}*0.7+0.1*\mathbf{1}$ $\mathbf{y}^{LS}=\mathbf{y}*0.7+0.15*\mathbf{1}$ KL Loss
Baseline $\mathbf{y}$ $\mathbf{y}$ Cross Entropy
Cross-Entropy: In a regular multi-classification task, cross-entropy is used to measure the difference between the model’s predicted results and the true labels. Assume there are classes, and each sample’s true label can be represented as a n-dimensional one-hot encoded vector p, where only one element is 1, indicating the corresponding class to which the sample belongs. And assume the model’s predicted result is also a n-dimensional probability distribution vector q, where each element represents the probability of the sample belonging to the corresponding class. The definition of the cross-entropy loss function H(p,q) is as follows:
$H(p,q)=-\sum_{i=1}^{n}p(x_{i})\log q(x_{i})$ (19)
The cross-entropy loss function calculates the negative logarithm of the product of the true label probability and the predicted probability for each class, summed over all classes. It penalizes the model for predicting low probabilities(Udhayakumar et al.,, 2019). Minimizing the cross-entropy loss function helps the model make more accurate predictions of the sample’s class.
KL Divergence: KL divergence measures the information difference between two probability distributions. In LS, it assesses the dissimilarity from the ground truth distribution to the predicted distribution. It emphasizes the discrepancy in information when transitioning from the true distribution to the predicted distribution. The formula is shown as follows:
$KL(p||q)=\sum_{i=1}^{n}p(x_{i})\log\left(\frac{p(x_{i})}{q(x_{i})}\right)$ (20)
Thus, we have:
$\begin{split}KL(p||q)=&-\sum_{i=1}^{n}p(x_{i})\log q(x_{i})\\ &+\sum_{i=1}^{n}p(x_{i})\log p(x_{i})\end{split}$ (21)
$KL(p||q)=H(p,q)-H(p)$ (22)
Although the formulas of cross-entropy and KL divergence differ only by the inclusion of the entropy of the true labels, their distinct implications and applications justify their separate usage(Chai et al.,, 2017). In LS, KL divergence is commonly preferred as an objective function due to the following reasons:
(I)Sensitivity to Distribution Shift: KL divergence is highly sensitive to shifts between distributions, making it suitable for capturing the accuracy of predicted distributions in reflecting the characteristics of the ground truth distribution, beyond just class labels.
(II)Handling Asymmetry: In certain scenarios, the true and predicted distributions may exhibit asymmetry, with varying degrees of dissimilarity. KL divergence handles this asymmetry better as it focuses on the information difference from the true distribution to the predicted distribution.
4.3 Results
The results produced by BERT, TEXTCNN and RoBERTa architectures are presented in Table 2. The accuracy curves are presented in Figure 4, Figure 5 and Figure 6 respectively, depicting the performance of the models at each epoch.
On the BERT architecture, LS models consistently outperform the baseline model across multiple datasets. Specifically, LS2 achieves the highest accuracy in most cases, closely followed by LS4 and LS5. This indicates that incorporating LS with varying degrees of LS can significantly improve classification accuracy, particularly for binary and ternary classification datasets.
Moving to the TextCNN architecture, LS models exhibit superior performance on ternary classification datasets compared to binary classification datasets. LS2 and LS4 consistently outperform the other models in terms of accuracy. This highlights the effectiveness of LS when implemented in the TextCNN architecture for enhancing classification outcomes. Although the training acceleration of LS models on the CNN architecture is slower, they have greater potential in the later epochs.
On the RoBERTa architecture, LS models demonstrate competitive performance, LS2 and LS5 achieving the highest accuracy in most cases. This suggests that LS can effectively leverage the strengths of the RoBERTa architecture to achieve improved classification accuracy.
The experimental results demonstrate that LS models, when integrated into different architectures, consistently outperform the baseline model. LS2, LS4, and LS5 consistently demonstrate superior accuracy, indicating the efficacy of LS in enhancing classification performance. These findings emphasize the potential of LS as a valuable approach for improving classification outcomes in various tasks and architectures.
The results emphasize the effectiveness of the LS approach in improving the accuracy of the models. The varying degrees of LS in the LS models provide a means to balance model sensitivity to noisy labels and mitigate overconfidence. This contributes to the models’ enhanced robustness and generalization capabilities, ultimately leading to improved performance in sentiment classification tasks.
Table 2: VALIDATION ACCURACY ON DIFFERENT ARCHITECTURES
@@ 4.3 Results
TextCNN LS2 0.8237 0.6852 0.7074 0.7761 0.8382 0.7234 0.7860 0.7455
LS3 0.8233 0.6801 0.7029 0.7740 0.8357 0.7245 0.7920 0.7540
LS4 0.8216 0.6792 0.7066 0.7771 0.8369 0.7183 0.7780 0.7455
LS5 0.8204 0.6843 0.7029 0.7730 0.8344 0.7338 0.7860 0.7444
Baseline 0.8141 0.6852 0.7049 0.7647 0.8319 0.7090 0.7920 0.7626
RoBERTa LS2 0.8915 0.8375 0.7946 0.8689 0.9031 0.8101 0.8680 0.8588
LS3 0.8961 0.8263 0.8036 0.8617 0.9070 0.8029 0.8660 0.8524
LS4 0.8957 0.8246 0.7999 0.8658 0.9108 0.8163 0.8720 0.8610
LS5 0.8978 0.8298 0.8024 0.8638 0.9045 0.8122 0.8780 0.8706
@@
Figure 4: Accuracy on the validation set at different epochs using BERT
Figure 5: Accuracy on the validation set at different epochs using TEXTCNN
Figure 6: Accuracy on the validation set at different epochs using RoBERTa
Figure 7: Visualization of Dimension Reduction of the BERT Model
5 Discussions
In this section, we attempt to analyze why text LS can achieve higher accuracy under the same data and parameter conditions.
Higher Accuracy: LS methods can achieve higher accuracy. The impact of label distribution can be profound, as it effectively addresses certain challenges in sentiment classification. Label distribution helps alleviate overconfidence in model predictions and reduces sensitivity to noisy labels, resulting in improved robustness and generalization performance.
@@ 5 Discussions
Acceleration of Convergence: We further analyzed the convergence performance of LS on the BERT, TextCNN, and RoBERTa architectures. It can be observed that in many cases, LS models implemented on BERT and RoBERTa exhibit significantly faster convergence compared to the baseline model, and can even achieve relatively high accuracy within the initial few epochs of training, as shown in Figure 4 and Figure 6. This also indicates that LS is capable of effectively saving computational resources, reducing the training time, and improving overall efficiency.
Robust Performance: The LS methods implemented on the three deep learning architectures consistently achieve stable and superior prediction performance on most of the datasets, demonstrating the robustness of the models. This indicates that the LS approach is effective in handling diverse data scenarios and can generalize well to unseen examples. Furthermore, the ability of the models to consistently outperform the baseline methods across multiple datasets highlights their reliability and suitability for real-world applications. The robust predictive performance of LS methods reinforces their value and reinforces their potential as a powerful tool in various domains requiring accurate and reliable predictions.
@@
Reduce Overfitting: LS and KL divergence can help reduce overfitting compared to cross-entropy for several reasons. Cross-entropy loss assigns high confidence to a single target class and penalizes all other classes. This can lead to overconfident predictions, especially when the training data is limited or imbalanced. In contrast, LS considers the entire label distribution, allowing the model to capture the relationships between classes and reduce the risk of overfitting to individual examples.
LS: LS further mitigates overfitting. The effects of LS can be reflected in the hidden layer of the Model. Figure 7 visualizes similarities and differences between data points by mapping high-dimensional features into two - dimensional spaces and depicts the last layer of BERT models with different smoothing levels. The data points in single-label model overlap, while model with smoothing labels can distinguish data points well. After applying the LS method, during the training phase, the loss doesn’t decrease too rapidly when the predictions are correct, and it doesn’t penalize too heavily when the predictions are wrong. This prevents the network from easily getting stuck in local optima and helps to mitigate overfitting to some extent. Additionally, in scenarios where the classification categories are closely related, the network’s predictions are not excessively absolute.
6 Conclusion
@@
In this paper, we delve into the application of LS in text sentiment classification, a key area of Natural Language Processing (NLP). Our experiments demonstrate that by fine-tuning the appropriate smoothing parameters, LS can achieve improved prediction accuracy across multiple deep learning architectures and various test datasets. However, given the inherent emotional bias in different texts, constructing more precise textual sentiment distribution labels to capture subtle variations in human emotions remains a challenging task. This highlights the need for further research and development in this area to fully harness the potential of LS in enhancing the performance and robustness of sentiment classification models.

### 23 Feb 2024 v2

@@ IV Experimental Results and Analysis
We construct four LS predictive models incorporating varying degrees of label smoothing and assess their performance compared to the baseline models on eight datasets using TextCNN, BERT, and RoBERTa architectures. In additional to evaluating the effectiveness of LS methods, we also conduct in-depth analyses of how LS benefits text sentiment classification.
IV-A Datasets
@@
We utilize publicly available datasets open-sourced in Kaggle and Huggingface. This experiment employs eight distinct sentiment analysis datasets, of which six are three-class and two are two-class. The diverse range of datasets enables us to evaluate the performance and generalizability of our models across various data types and scenarios.
@@
Twitter Financial News Sentiment (TFNS) was collected using the Twitter API. It holds 11,932 documents annotated with 3 labels (’Positive’, ’Neutral’, ’Negatve’) and comprises annotated finance-related tweets.
@@
Kaggle Financial Sentiment (KFS) is a dataset of financial reviews annotated with 3 labels (’Positive’, ’Neutral’, ’Negatve’). These reviews cover topics such as stocks, investments, market analysis, corporate performance, technology, and real estate.
@@
Tweet Sentiment Extraction (TSE) is a three-class dataset from a Kaggle contest. It has 27481 samples of a tweet and a sentiment label (’Positive’, ’Neutral’, or ’Negatve’).
@@
Auditor Sentiment (AS) gathers the auditor evaluations into one dataset. It contains thousands of sentences from English financial news, grouped into three categories by emotion (’Positive’, ’Neutral’, ’Negatve’).
@@
Financial Phrasebank (FP) consists of 4840 sentences from English language financial news categorised by sentiment. The dataset is divided into three classes (’Positive’, ’Neutral’, ’Negatve’) by agreement rate of 5-8 annotators.
@@
Chatgpt Sentiment Analysis (CSA) contains 10,000 pieces of data. The dataset gathers tweets about Chatgpt and Reflects people’s perception of it. Emotions are divided into three categories (’Positive’, ’Neutral’, ’Negatve’).
@@
Rotten Tomatoes Reviews (RTR) consists of 10,662 processed sentences from Rotten Tomatoes movie reviews. It’s a balanced two-class dataset with 5,331 positive and 5,331 negative.
@@
Sentiment140 (Sent140) is a widely used sentiment analysis dataset created by researchers at Stanford University. It contains 1,600,000 tweets extracted using the twitter api. The tweets have been annotated to two classes (’Positive’, ’Negatve’).
IV-B Model Configuration
@@
Models: Table I presents the configuration of the models and loss functions used in the experiment. The LS models (LS1–LS4) employ different smoothing levels, with the smoothing parameter $\lambda$ increasing from 0.01 (LS1) to 0.1 (LS4). The original one-hot hard labels are used in the baseline methods, while for LS1 to LS4, the soft labels are utilized with different smoothing parameters. By gradually increasing the smoothing parameter from LS1 to LS4, we explore how label smoothing affects the model’s accuracy and stability. All LS models utilize the KL divergence as the loss function. In contrast, the baseline models employ the cross entropy loss.
@@
TABLE I: Model smoothing levels and loss functions. The baseline methods take the original one-hot hard label, $\mathbf{y}$. We vary the smoothing parameter $\lambda$ to obtain four sets of soft labels $\mathbf{y}^{LS}$.
@@
Model Smoothing para.(three/two-class) Smoothed label (three-class) Smoothed Label (two-class) Loss
Baseline $\lambda=0/0$ $\mathbf{y}$ $\mathbf{y}$ Cross Entropy
LS1 $\lambda=0.01/0.01$ $\mathbf{y}^{LS}=\mathbf{y}*0.97+0.01*\mathbf{1}$ $\mathbf{y}^{LS}=\mathbf{y}*0.98+0.01*\mathbf{1}$ KL Loss
LS2 $\lambda=0.025/0.05$ $\mathbf{y}^{LS}=\mathbf{y}*0.925+0.025*\mathbf{1}$ $\mathbf{y}^{LS}=\mathbf{y}*0.9+0.05*\mathbf{1}$ KL Loss
LS3 $\lambda=0.05/0.1$ $\mathbf{y}^{LS}=\mathbf{y}*0.85+0.05*\mathbf{1}$ $\mathbf{y}^{LS}=\mathbf{y}*0.8+0.1*\mathbf{1}$ KL Loss
LS4 $\lambda=0.1/0.15$ $\mathbf{y}^{LS}=\mathbf{y}*0.7+0.1*\mathbf{1}$ $\mathbf{y}^{LS}=\mathbf{y}*0.7+0.15*\mathbf{1}$ KL Loss
TABLE II: The test accuracy of different classifiers on eight sentiment classification datasets with varying smoothing parameters.
@@
TextCNN LS1 0.8237 0.6852 0.7074 0.7761 0.8382 0.7234 0.7860 0.7455
LS2 0.8233 0.6801 0.7029 0.7740 0.8357 0.7245 0.7920 0.7540
LS3 0.8216 0.6792 0.7066 0.7771 0.8369 0.7183 0.7780 0.7455
LS4 0.8204 0.6843 0.7029 0.7730 0.8344 0.7338 0.7860 0.7444
Baseline 0.8141 0.6851 0.7049 0.7647 0.8319 0.7090 0.7918 0.7626
RoBERTa LS1 0.8915 0.8375 0.7946 0.8689 0.9031 0.8101 0.8680 0.8588
LS2 0.8961 0.8263 0.8036 0.8617 0.9070 0.8029 0.8660 0.8524
LS3 0.8957 0.8246 0.7999 0.8658 0.9108 0.8163 0.8720 0.8610
LS4 0.8978 0.8298 0.8024 0.8638 0.9045 0.8122 0.8780 0.8706
@@
Figure 2: Accuracy on the validation set at different epochs using BERT with varying smoothing parameters.
IV-C Results
Performance: We employ accuracy as the main metrics to evaluate the performance of methods on eight datasets. The results produced by BERT, TextCNN and RoBERTa architectures are presented in Table II. TextCNN models are trained from scratch with random initializations, while BERT and RoBERTa models are finetuned on sentiment analysis datasets with pre-trained weights. On the BERT architecture, LS models consistently outperform the baseline model across multiple datasets. Specifically, LS1 achieves the highest accuracy in most cases, closely followed by LS3 and LS4. This indicates that incorporating LS with varying degrees of LS can significantly improve classification accuracy, particularly for binary and ternary classification datasets.
Moving to the TextCNN architecture, LS models exhibit superior performance on three-category classification datasets compared to binary classification datasets. LS1 and LS3 consistently outperform the other models in terms of accuracy. This highlights the effectiveness of LS when implemented in the TextCNN architecture for enhancing classification outcomes.
On the RoBERTa architecture, LS models demonstrate competitive performance, LS1 and LS4 achieving the highest accuracy in most cases. This suggests that LS can effectively leverage the strengths of the RoBERTa architecture to achieve improved classification accuracy.
The experimental results demonstrate that LS models, when integrated into different architectures, consistently outperform the baseline model. LS1, LS3, and LS4 consistently demonstrate superior accuracy, indicating the efficacy of LS in enhancing classification performance.
(a) t-SNE plot from BERT baseline
(b) t-SNE plot from BERT with LS1
(c) t-SNE plot from BERT with LS2
(d) t-SNE plot from BERT with LS3
Figure 3: t-SNE plots from BERT baseline versus label smoothing methods.
IV-D Analysis
In this section, we attempt to analyze why and how LS benefits the text sentiment classification.
Higher Accuracy: As shown in Table II, LS methods achieve higher accuracy than baseline models. The impact of label distribution can be profound, as it effectively addresses certain challenges in sentiment classification. Label distribution helps alleviate overconfidence in model predictions and reduces sensitivity to noisy labels, resulting in improved robustness and generalization performance.
@@ IV-D Analysis
Acceleration of Convergence: We further analyzed the convergence performance of deep learning architectures during the training of sentiment classifiers. The accuracy curves of BERT are presented in Figure 2, depicting the performance of the models on the validation sets at each training epoch. It can be observed that LS models implemented on BERT exhibit significantly faster convergence compared to the baseline model, and can even achieve relatively high accuracy within the initial few epochs of training. This also indicates that LS is capable of effectively saving computational resources, reducing the training time, and improving overall efficiency.
@@
Reduce Overfitting: LS and KL divergence can help reduce overfitting compared to cross-entropy for several reasons. Cross-entropy loss assigns high confidence to a single target class and penalizes all other classes. This can lead to overconfident predictions, especially when the training data is limited or imbalanced. In contrast, LS considers the entire label distribution, allowing the model to capture the relationships between classes and reduce the risk of overfitting to individual examples. After applying the LS method, during the training phase, the loss does not decrease too rapidly when the predictions are correct, and it does not penalize too heavily when the predictions are wrong. This prevents the network from easily getting stuck in local optima and helps to mitigate overfitting to some extent. Additionally, in scenarios where the classification categories are closely related, the network’s predictions are not excessively absolute.
Representation Learning: LS can produce better hidden representations for texts than baseline methods. The effects of LS can be reflected in the hidden layer of the Model. Figure 3 visualizes similarities and differences between data points by mapping high-dimensional features into two-dimensional spaces and depicts the last layer of BERT models with different smoothing levels. The data points from baseline BERT model overlap as shown in Fig. 2(a), while datapoints from LS methods are linearly separable as shown in Fig 2(b)2(d).
V Conclusion
@@
In this paper, we delve into the application of LS in text sentiment classification, a key problem of natural language processing. Our experiments demonstrate that by tuning the smoothing parameters, LS can achieve improved prediction accuracy across multiple deep learning architectures and various test datasets. We conduct fine-grained analysis to check how LS benefits the task. Our main findings are: 1.) LS can boost the convergence of deep models in both training from scratch and fine-tuning modes; 2.) LS can produce hidden representations for texts that make the data of different labels more easily separable. However, given the inherent emotional bias in different texts, constructing more precise textual sentiment distribution labels to capture subtle variations in human emotions remains a challenging task. This highlights the need for further research and development in this area to fully harness the potential of LS in enhancing the performance and robustness of sentiment classification models.